OpenStudy (anonymous):

Given A(9, 11), B(3, -4), C(6, 9), and D(-7, -7), then the midpoint of which line is located in Quadrant III? AB BC CD BD If A(5, 5), B(3, 3), C(6, 8), and D(-3, 4), then which segment is the longest? AB BC CD BD

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

mid point of AB is: \[(\frac{ 9+3 }{ 2 },\frac{ 11-4 }{ 2 })=(6,7/2)\] mid point of CD, is:

OpenStudy (michele_laino):

\[\left( \frac{ 6-7 }{ 2 } ,\frac{ 9-7 }{ 2 }\right)=(-1/2,1)\]

OpenStudy (michele_laino):